prove that the function u = 2x (1-y) is harmonic


1
Expert's answer
2021-08-31T16:22:53-0400
"u=2x(1-y)"

"\dfrac{\partial u}{\partial x}=2(1-y)"

"\dfrac{\partial^2 u}{\partial x^2}=0"

"\dfrac{\partial u}{\partial y}=-2x"

"\dfrac{\partial^2 u}{\partial y^2}=0"

"\nabla^2u(x,y)=\dfrac{\partial^2 u}{\partial x^2}+\dfrac{\partial^2 u}{\partial y^2}"

"=0+0=0"

Then the function "u = 2x (1-y)" is harmonic.